OpenStudy (aonz):

Help please. I just need someone to explain to be something about Inequations and inequalities(Parabolas). Please know what your talking about btw :)

OpenStudy (aonz):

There is a type of question i always get wrong and i have no idea why,

OpenStudy (anonymous):

You forgot to reject the answer \(x=0\), since it doesn't work with the original expression (division by zero is undefined). Your method is correct, though.

OpenStudy (aonz):

so x can never equal 0?

OpenStudy (aonz):

i feel stupid asking this but why cant x=0?

OpenStudy (anonymous):

No, not never. Just for this problem, since \(\dfrac{1}{x}\) is in the original expression.

OpenStudy (aonz):

ok i drew a parabola and i got \[-4\le x \le-\frac{ 5 }{ 2 }\]

OpenStudy (aonz):

book says \[-4<x \le \frac{ 5 }{ 2 }\]

OpenStudy (aonz):

original question is